# Enhancing Physics-Informed Neural Networks Through Feature Engineering

### v2

@@ Abstract
Physics-Informed Neural Networks (PINNs) seek to solve partial differential equations (PDEs) with deep learning. Mainstream approaches that deploy fully-connected multi-layer deep learning architectures require prolonged training to achieve moderate accuracy, while recent work on feature engineering allows higher accuracy and faster convergence. This paper introduces SAFE-NET, a Single-layered Adaptive Feature Engineering NETwork that improves errors with far fewer parameters than baseline feature engineering methods. SAFE-NET returns to basic ideas in machine learning, using Fourier features, a simplified single hidden layer network architecture, and an effective optimizer that improves the conditioning of the PINN optimization problem. Numerical results show that SAFE-NET converges faster and typically outperforms deeper networks and more complex architectures. It consistently uses fewer parameters — on average, 53% fewer than the competing feature engineering methods and $\boldsymbol{\times}$ fewer than state-of-the-art large-scale architectures — while achieving comparable accuracy in less than 30% of the training epochs. Moreover, each SAFE-NET epoch is 95% faster than those of competing feature engineering approaches. These findings challenge the prevailing belief that modern PINNs effectively learn relevant features and highlight the efficiency gains possible through feature engineering.
@@ 1 Introduction
Partial Differential Equations (PDEs) underpin scientific modeling but remain notoriously challenging to solve. Classical numerical methods struggle with high dimensionality and nonlinearity, while analytical solutions are rare. Physics-Informed Neural Networks (PINNs) (Raissi et al., 2019a; Karniadakis et al., 2021) have emerged as a promising alternative, leveraging neural networks to approximate PDE solutions through residual minimization. They aim to solve PDE systems of the form
$D[u(x),x]$ $=0,$ $\quad x\in\Omega$ (1)
$B[u(x),x]$ $=0,$ $\quad x\in\partial\Omega$
$I[u(x),x]$ $=0,$ $\quad x\in\Omega$
@@
Where $D$ represents the differential operator defining the PDE, $B$ represents the boundary conditions. $I$ represents the initial conditions, important for time-dependent problems, and $\Omega\subseteq\mathbb{R}^{d}$ is the domain of the PDE. By avoiding mesh generation, PINNs offer flexibility for forward/inverse problems and high-dimensional settings.
@@
Despite their potential, PINNs face a fundamental challenge—they are difficult to train (Krishnapriyan et al., 2021; Rathore et al., 2024). Research shows that the differential operator in the residual loss induces ill-conditioning (De Ryck et al., 2023; Rathore et al., 2024), leading to a poor optimization landscape and slow convergence for popular first-order optimizers such as Adam (Kingma & Ba, 2014). Thus, successful PINN training can be time-consuming and fiddly, limiting the use of PINNs.
@@
Recent work has developed several strategies to improve PINN training, including feature engineering. Feature engineering endows the network with additional features that better capture the inductive bias of the learning task. A range of feature engineering approaches, from Fourier features (RFF) (Wang et al., 2020) to radial basis function features (RBF) (Zeng et al., 2024a), have been proposed.This focus on feature engineering represents a departure from the dominant trend in modern deep learning, where end-to-end learning has largely replaced hand-crafted features. However, the scientific computing domain presents unique challenges that may favor explicit feature design. Unlike image recognition or natural language processing, where optimal feature representations are unknown a priori, PDE solutions have well-understood mathematical structures rooted in Fourier analysis, differential geometry, and physics principles. This domain knowledge provides strong theoretical guidance for feature design that is often absent in other ML applications. Recent work in domain-specific architectures—from Graph Neural Networks for relational data (Bronstein et al., 2017) to Transformer attention mechanisms for sequences (Vaswani et al., 2023)—suggests that incorporating structural knowledge can outperform generic architectures.
@@
    Domain knowledge features encode physical priors, derived from boundary conditions or initial conditions.
@@
SAFE-NET does not directly target the optimization landscape like prior feature engineering approaches such as RFFs or RBFs. Instead, it seeks to improve the inductive bias of the PINN by augmenting the initial data representation with well-conditioned Fourier features and domain knowledge. These design choices makes SAFE-NET particularly effective for PDEs without shocks or discontinuities, where Fourier bases work best. Interestingly, this inductive bias also provides an implicit preconditioning effect that leads to a better-conditioned optimization problem and facilitates training.
@@
  • SAFE-NET achieves runtime performance comparable to non-feature engineering methods with very low parameter count, while other feature engineering approaches are slower (Table 1 and Figure 4).
  • Experiments across a wide variety of PDEs and baseline methods (Table 2) show that on average, SAFE-NET yields the best or comparable performance.
Deep learning has been seen as a promising tool for solving PDEs. However, our work with SAFE-NET shows that traditional techniques like feature engineering can achieve lower error rates and faster training times than conventional multi-layer networks. While recent large-scale approaches such as PINNsFormer (Zhao et al., 2024), PirateNets (Wang et al., 2024), and operator learning approaches such as FNO (Li et al., 2020b) and its variations may achieve superior accuracy in certain PDEs through complex architectures, these approaches are extremely expensive as they use very large networks: e.g., $454$k parameters for PINNsFormer (Zhao et al., 2023), $\geq 720$k parameters for PirateNets (Wang et al., 2024), and approximately $1$M parameters for FNO and similar approaches. This paper shows that for many PDE tasks, simple feature engineering can perform as well or better than complex architectures, challenging the preconception that deeper networks effectively learn important problem features.
@@ 2.1 Fourier-Based Feature Engineering
Fourier feature mappings leverage the spectral properties of PDE solutions to enhance high-frequency learning. They aim to address spectral bias — the tendency of neural networks to favor low-frequency functions — by transforming input coordinates into a more expressive representation. The most prominent approach, RFF (Wang et al., 2021b), uses the feature mapping
@@
with fixed Gaussian weights $B\in\mathbb{R}^{m\times d}$ drawn from $\mathcal{N}(0,\sigma^{2})$. With both cosine and sine terms, this mapping projects inputs into a high-dimensional space where periodic and high-frequency patterns are more easily captured. Theoretical insights from (Tancik et al., 2020b) show that Fourier features help neural networks learn high-frequency functions in low-dimensional domains, which are typical in PDE applications. However, RFF’s reliance on fixed, random frequencies limits its adaptation to PDE-specific spectral properties. Moreover, the Gaussian initialization of $B$ may not align with the dominant frequencies of the solution, leading to reduced performance.
Another baseline method we study, RBA-PINN from (Anagnostopoulos et al., 2023), also incorporates Fourier feature embeddings for enforcing periodic boundary conditions in some of its variants. While the main contribution of RBA-PINN is the residual-based attention weighting scheme, their ablation studies demonstrate that Fourier features provide the most significant performance improvement among all components tested. Specifically, for the Allen-Cahn equation, RBA alone achieves $3.16\times 10^{-3}$ relative $L^{2}$ error, while RBA+Fourier improves this by two orders of magnitude. The authors explicitly acknowledge that the coupling of the RBA scheme along with the Fourier feature embedding is the most important component to achieving low L2RE. We therefore include RBA-PINN as a feature engineering baseline (for periodic boundary conditions), recognizing that while RBA’s core innovation lies in adaptive weighting, its best-performing configuration on PDEs with periodic boundaries relies heavily on Fourier feature engineering for its effectiveness.
@@ 2.2 Non-Fourier Approaches
For PDEs with sharp gradients or discontinuities, Fourier features may struggle due to the Gibbs phenomenon; see (Zeng et al., 2024b). The Burgers PDE, with its sharp discontinuity at $x=0$, is well-suited to observe this behavior. Comparing numerical results for the Burgers PDE across methods in Table 3 show that RFF performs comparably to methods without feature engineering but underperforms RBF by over an order of magnitude.
(Zeng et al., 2024a) addresses the Gibbs phenomenon by using Radial Basis Functions (RBFs) as
$\phi_{\text{RBF}}(x)=\exp\!\Bigl{(}-\tfrac{\lvert x-c\rvert^{2}}{2\sigma^{2}}\Bigr{)},$ (3)
where $c$ denotes the center and $\sigma$ controls the kernel width. RBF expansions can better approximate local, abrupt changes but are computationally intensive due to kernel regression requirements and less suited for periodic or high-frequency PDE solutions (Tancik et al., 2020b). The RBF feature mapping function is defined as:
$\Phi(x)=\frac{\sum_{i}^{m}w_{i}\phi_{\text{RBF}}(|x-c_{i}|)}{\sum_{i}^{m}\phi_{\text{RBF}}(|x-c_{i}|)}$ (4)
where $c\in\mathbb{R}^{n\times m}$ are the centers of the RBFs (trainable parameters) and $w$ is the weight matrix for the feature mapping layer.
Polynomial Enhancement. A key limitation of pure RBF approaches is their difficulty in representing global trends and polynomial components that frequently appear in PDE solutions. To address this, RBF-P (RBF with Polynomials) from (Zeng et al., 2024b) augments the RBF feature set with polynomial terms, so the feature mapping function is adjusted to
@@
where P is the polynomial function. In the feature mapping layer, it can be represented as $\begin{bmatrix}f_{1}\\ \vdots\\ f_{N}\end{bmatrix}$ equal to
$\begin{bmatrix}\phi_{\text{RBF}}(r_{1}^{1})&\cdots&\phi_{\text{RBF}}(r_{1}^{m})&\mid&1&x_{1}&x^{k}\\ \vdots&\ddots&\vdots&\mid&\vdots&\vdots&\vdots\\ \phi_{\text{RBF}}(r_{N}^{1})&\cdots&\phi_{\text{RBF}}(r_{N}^{m})&\mid&1&x_{N}&x_{N}^{k}\end{bmatrix}\begin{bmatrix}W^{m}\\ W^{k}\end{bmatrix}$
@@ 3.1 Theoretical Background
$f(\mathbf{x})\approx\sum_{\mathbf{\kappa}\text{ dominant}}\left(A_{\mathbf{\kappa}}\cos(2\pi\mathbf{\kappa}\cdot\mathbf{x})+B_{\mathbf{\kappa}}\sin(2\pi\mathbf{\kappa}\cdot\mathbf{x})\right),$
@@
Figure 2: Diagram showing how SAFE-NET works for a 1D time-dependent PDE as an example. The Feature Generator Module has trainable frequencies and coefficient for more effective feature selection.
@@
[…]
@@
[…]
This expansion produces four real-valued basis functions per frequency pair $(\omega_{x},\lambda_{t})$ as
@@ 3.2 SAFE-NET
Motivated by the considerations of Section 3.1, SAFE-NET implements the parametric basis in equations (5)-(8) through learnable frequencies $\{\omega_{x}^{(i)},\lambda_{t}^{(i)}\}_{i=1}^{N}$ and amplitudes $\{a^{(i)},b^{(i)},c^{(i)},d^{(i)}\}_{i=1}^{N}$ to estimate
$u_{\theta}(x,t)$ $=\sum_{i=1}^{N}\big{[}a^{(i)}\phi_{1}^{\omega_{x},\lambda_{t}}(x,t)+b^{(i)}\phi_{2}^{\omega_{x},\lambda_{t}}(x,t)$
$\quad+c^{(i)}\phi_{3}^{\omega_{x},\lambda_{t}}(x,t)+d^{(i)}\phi_{4}^{\omega_{x},\lambda_{t}}(x,t)\big{]}$ (9)
where $\theta=\{\omega_{x}^{(i)},\lambda_{t}^{(i)},a^{(i)},b^{(i)},c^{(i)},d^{(i)}\}$ are trainable parameters. The explicit cross-frequency terms in equation (9) capture the tensor product structure of the 2D Fourier basis.
@@
Domain Knowledge features for the PDEs considered in this paper appear in Appendix A.1. As shown in Figure 2, depending on the availability of domain information, these features are concatenated with the Fourier basis terms before normalization and linear projection. SAFE-NET uses only one hidden layer: given an input $(\bm{x},t)\in\Omega\times\mathbb{R}$, the SAFE-NET network computes
@@
with parameters $\theta=(w_{\textup{SAFE-NET}},W_{1},b_{1},w_{2},b_{2})$, nonlinearity $\sigma=\tanh(\cdot),$ and learnable feature mapping $\phi(w_{\textup{SAFE-NET}},(\bm{x},t))$. Complete architectural specifications, parameter initialization schemes, and training procedures for SAFE-NET are provided in Appendix A.1. To ensure robust experimental validation, we conducted comprehensive ablation studies examining key design choices including sensitivity to the number of Fourier features (see Figure 9), activation function selection for the single hidden layer, and sensitivity to our initialization strategy for Fourier coefficients and frequencies. These ablation results, presented Appendix C.1, explaining the choice of each component.
By combining the generality of Fourier features with known solution characteristics, SAFE-NET accelerates convergence. Figure 3 compares SAFE-NET with the best competing PINN architecture for each PDE in Table 2.

### v3

@@ Abstract
[…]
@@ 1 Introduction
Partial Differential Equations (PDEs) underpin scientific modeling but remain notoriously challenging to solve. Classical numerical methods struggle with high dimensionality and nonlinearity, while analytical solutions are rare. Traditional numerical methods are supported by software libraries such as FEniCS  (Alnaes2015), deal.II (Arndt_2021) and PETSc (petsc-web-page), but still considerable domain expertise is required for effective implementation. More recently, LLM-based approaches have been explored for automating PDE solver generation, combining symbolic reasoning and code synthesis to produce solvers (li2025codepdeinferenceframeworkllmdriven; soroco2025pdecontrollerllmsautoformalizationreasoning; fazliani2025pdesharppdesolverhybrids). Such approaches might struggle when the target PDE lacks prior numerical treatment or is underrepresented in the training data, limiting their generality. Physics-Informed Neural Networks (PINNs) (raissi2019unified; karniadakis2021physicsinformed) are a promising alternative, leveraging neural networks to approximate PDE solutions through residual minimization. They aim to solve PDE systems of the form
$D[u(\mathbf{x}),\mathbf{x}]$ $=0,$ $\quad\mathbf{x}\in\Omega$ (1)
$B[u(\mathbf{x}),\mathbf{x}]$ $=0,$ $\quad\mathbf{x}\in\partial\Omega$
@@
where $D$ represents the differential operator defining the PDE, $B$ represents a general boundary condition operator, and $\Omega\subseteq\mathbb{R}^{d}$ is the domain of the PDE. By avoiding mesh generation, PINNs offer flexibility for forward/inverse problems and high-dimensional settings.
@@
Despite their potential, PINNs face a fundamental challenge—they are difficult to train (krishnapriyan2021characterizing; rathore2024challenges). Research shows that the differential operator in the residual loss induces ill-conditioning (de2023operator; rathore2024challenges), leading to a poor optimization landscape and slow convergence for popular first-order optimizers such as Adam (kingma2014adam). Thus, successful PINN training can be time-consuming and fiddly, limiting the use of PINNs.
@@
Recent work has developed several strategies to improve PINN training, including feature engineering. Feature engineering endows the network with additional features that better capture the inductive bias of the learning task. A range of feature engineering approaches, from Fourier features (RFF) (wang2020eigenvector) to radial basis function features (RBF) (zeng2024featuremappingphysicsinformedneural), have been proposed.This focus on feature engineering represents a departure from the dominant trend in modern deep learning, where end-to-end learning has largely replaced hand-crafted features. However, the scientific computing domain presents unique challenges that may favor explicit feature design. Unlike image recognition or natural language processing, where optimal feature representations are unknown a priori, PDE solutions have well-understood mathematical structures rooted in Fourier analysis, differential geometry, and physics principles. This domain knowledge provides strong theoretical guidance for feature design that is often absent in other ML applications. Recent work in domain-specific architectures—from Graph Neural Networks for relational data Bronstein_2017 to Transformer attention mechanisms for sequences vaswani2017attention—suggests that incorporating structural knowledge can outperform generic architectures.
@@
    Domain knowledge features derived from initial and boundary conditions (if available), providing added benefits in boosting accuracy.
@@
SAFE-NET does not directly target the optimization landscape like prior feature engineering approaches such as RFFs or RBFs. Instead, it seeks to improve the inductive bias of the PINN by augmenting the initial data representation with well-conditioned Fourier features and domain knowledge. These design choices make SAFE-NET particularly effective for PDEs without shocks or discontinuities, where Fourier bases work best. Interestingly, this inductive bias also provides an implicit preconditioning effect that leads to a better-conditioned optimization problem and facilitates training.
@@
  • SAFE-NET achieves runtime performance comparable to non-feature engineering methods with very low parameter count, while other feature engineering approaches are slower (Table˜1 and Figure˜4).
  • Experiments across a wide variety of PDEs, including (i) linear, mildly nonlinear, and highly nonlinear dynamics, (ii) time‑dependent and steady‑state problems, and (iii) PDEs with predominantly smooth behavior as well as discontinuous shock-forming PDEs, in dimensions $\leq 3$ and baseline methods (Table˜2) show that on average, SAFE-NET yields the best or comparable performance.
Deep learning has been seen as a promising tool for solving PDEs. However, our work with SAFE-NET shows that traditional techniques like feature engineering can achieve lower error rates and faster training times than conventional multi-layer networks, particularly suited to PDEs whose solutions are sufficiently smooth. While recent large-scale approaches such as PINNsFormer zhao2024pinnsformertransformerbasedframeworkphysicsinformed, PirateNets wang2024piratenets, and operator learning approaches such as FNO li2020fno and its variations may achieve superior accuracy in certain PDEs through complex architectures, these approaches are extremely expensive as they use very large networks: e.g., 454454k parameters for PINNsFormer zhao2023pinnsformer, $\geq 720$k parameters for PirateNets wang2024piratenets, and approximately 11M parameters for FNO and similar approaches. This paper shows that for many PDE tasks, simple feature engineering can perform as well or better than complex architectures, challenging the preconception that deeper networks effectively learn important problem features.
@@ 2.1 Fourier-Based Feature Engineering
Fourier feature mappings leverage the spectral properties of PDE solutions to enhance high-frequency learning. They aim to address spectral bias — the tendency of neural networks to favor low-frequency functions — by transforming input coordinates into a more expressive representation. The most prominent approach, RFF wang2021eigenvector, uses the feature mapping
@@
with fixed Gaussian weights $B\in\mathbb{R}^{m\times d}$ drawn from $\mathcal{N}(0,\sigma^{2})$. With both cosine and sine terms, this mapping projects inputs into a high-dimensional space where periodic and high-frequency patterns are more easily captured. Theoretical insights from tancik2020fourierfeaturesletnetworks show that Fourier features help neural networks learn high-frequency functions in low-dimensional domains, which are typical in PDE applications. However, RFF’s reliance on fixed, random frequencies limits its adaptation to PDE-specific spectral properties. Moreover, the Gaussian initialization of $B$ may not align with the dominant frequencies of the solution, leading to reduced performance.
Another baseline method we study, RBA-PINN from anagnostopoulos2023residualbasedattentionconnectioninformation, also incorporates Fourier feature embeddings for enforcing periodic boundary conditions in some of its variants. While the main contribution of RBA-PINN is the residual-based attention weighting scheme, their ablation studies demonstrate that Fourier features provide the most significant performance improvement among all components tested. Specifically, for the Allen-Cahn equation, RBA alone achieves $3.16\times 10^{-3}$ relative $L^{2}$ error, while RBA+Fourier improves this by two orders of magnitude. The authors explicitly acknowledge that the coupling of the RBA scheme along with the Fourier feature embedding is the most important component to achieving low L2RE. We therefore include RBA-PINN as a feature engineering baseline (for periodic boundary conditions), recognizing that while RBA’s core innovation lies in adaptive weighting, its best-performing configuration on PDEs with periodic boundaries relies heavily on Fourier feature engineering for its effectiveness.
@@ 2.2 Non-Fourier Approaches
For PDEs with sharp gradients or discontinuities, Fourier features may struggle due to the Gibbs phenomenon; see zeng2024rbfpinnnonfourierpositionalembedding. The Burgers PDE, with its sharp discontinuity at $x=0$, is well-suited to observe this behavior. Comparing numerical results for the Burgers PDE across methods in Table˜3 show that RFF performs comparably to methods without feature engineering but underperforms RBF by over an order of magnitude.
zeng2024featuremappingphysicsinformedneural addresses the Gibbs phenomenon by using Radial Basis Functions (RBFs) as
$\phi_{\text{RBF}}(x;c_{i},\sigma)=\exp\!\left(-\frac{\|x-c_{i}\|^{2}}{2\sigma^{2}}\right),$ (3)
where $c_{i}$ denotes the center and $\sigma$ controls the kernel width. RBF expansions can better approximate local, abrupt changes but are computationally intensive due to kernel regression requirements and less suited for periodic or high-frequency PDE solutions tancik2020fourierfeaturesletnetworks. The RBF feature mapping function is defined as:
$\Phi(x)=\frac{\sum_{i=1}^{m}w_{i}\,\phi_{\text{RBF}}(x;c_{i},\sigma)}{\sum_{i=1}^{m}\phi_{\text{RBF}}(x;c_{i},\sigma)},$ (4)
where $\{c_{i}\}$’s are the centers of the RBFs (trainable parameters) and $\{w_{i}\}$’s are weights for the feature mapping layer.
Polynomial Enhancement. A key limitation of pure RBF approaches is their difficulty in representing global trends and polynomial components that frequently appear in PDE solutions. To address this, RBF-P (RBF with Polynomials) from zeng2024rbfpinnnonfourierpositionalembedding augments the RBF feature set with polynomial terms, so the feature mapping function is adjusted to
@@
where P is the polynomial function. In the feature mapping layer, it can be represented as
$\begin{bmatrix}f_{1}\\ \vdots\\ f_{N}\end{bmatrix}=\begin{bmatrix}\phi_{\text{RBF}}(r_{1}^{1})&\cdots&\phi_{\text{RBF}}(r_{1}^{m})&\mid&1&x_{1}&x^{k}\\ \vdots&\ddots&\vdots&\mid&\vdots&\vdots&\vdots\\ \phi_{\text{RBF}}(r_{N}^{1})&\cdots&\phi_{\text{RBF}}(r_{N}^{m})&\mid&1&x_{N}&x_{N}^{k}\end{bmatrix}\begin{bmatrix}W^{m}\\ W^{k}\end{bmatrix}$
@@ 3.1 Theoretical Background
[…]
@@
Figure 2: Diagram showing how SAFE-NET works for a 1D time-dependent PDE as an example. The Feature Generator Module has trainable frequencies and coefficients for more effective feature selection.
@@
$\hat{u}(\omega_{x},\lambda_{t})$ $=\iint_{\mathbb{R}^{2}}u(x,t)e^{-2\pi i(\omega_{x}x+\lambda_{t}t)}dxdt,$
$u(x,t)$ $=\iint_{\mathbb{R}^{2}}\hat{u}(\omega_{x},\lambda_{t})e^{2\pi i(\omega_{x}x+\lambda_{t}t)}d\omega_{x}d\lambda_{t},$
@@
$=\begin{aligned} \big[\cos(2\pi\omega_{x}x)+i\sin(2\pi\omega_{x}x)\big]\\ \otimes\big[\cos(2\pi\lambda_{t}t)+i\sin(2\pi\lambda_{t}t)\big].\end{aligned}$
For notational simplicity we absorb $2\pi$ into $\omega_{x},\lambda_{t}$ in what follows. This expansion produces four real-valued basis functions per frequency pair $(\omega_{x},\lambda_{t})$ as
@@ 3.2 SAFE-NET
Motivated by the considerations of Section˜3.1, SAFE-NET implements the parametric basis in equations (5)-(8) through learnable frequencies $\{\omega_{x}^{(\ell)},\lambda_{t}^{(\ell)}\}_{i=1}^{N}$ and amplitudes $\{a^{(\ell)},b^{(\ell)},c^{(\ell)},d^{(\ell)}\}_{i=1}^{N}$ to estimate
$u_{\theta}(x,t)$ $=\sum_{\ell=1}^{N}\big[a^{(\ell)}\phi_{1}^{\omega_{x}^{(\ell)},\lambda_{t}^{(\ell)}}(x,t)+b^{(\ell)}\phi_{2}^{\omega_{x}^{(\ell)},\lambda_{t}^{(\ell)}}(x,t)$
$\quad+c^{(\ell)}\phi_{3}^{\omega_{x}^{(\ell)},\lambda_{t}^{(\ell)}}(x,t)+d^{(\ell)}\phi_{4}^{\omega_{x}^{(\ell)},\lambda_{t}^{(\ell)}}(x,t)\big]$ (9)
where $\theta=\{\omega_{x}^{(\ell)},\lambda_{t}^{(\ell)},a^{(\ell)},b^{(\ell)},c^{(\ell)},d^{(\ell)}\}$ are trainable parameters. The explicit cross-frequency terms in equation (9) capture the tensor product structure of the 2D Fourier basis.
@@
  • For Boundary Conditions: If homogeneous Dirichlet conditions $u(0,t)=u(L,t)=0$ are present, features satisfying these conditions naturally (e.g., $\sin(\frac{n\pi x}{L})$) are prioritized by SAFE-NET’s frequency initialization (Appendix A.1.4).
Domain Knowledge features for the PDEs considered in this paper appear in Appendix A.1. As shown in Figure 2, depending on the availability of domain information, these features are concatenated with the Fourier basis terms before normalization and linear projection. Ablation studies on the role of these extra added features (see Appendix C.1.6) show that they result in improvements varying across different PDE tasks. Overall, they provide more accuracy boosting when the initial and boundary conditions carry more information of the final solution while being a lightweight supplementary component of SAFE-NET. SAFE-NET uses only one hidden layer: given an input $(\bm{x},t)\in\Omega\times\mathbb{R}$, the SAFE-NET network computes
@@
with parameters $\theta=(w_{\textup{SAFE-NET}},W_{1},b_{1},w_{2},b_{2})$, nonlinearity $\sigma=\tanh(\cdot),$ and learnable feature mapping $\phi(w_{\textup{SAFE-NET}},(\bm{x},t))$. Complete architectural specifications, parameter initialization schemes, and training procedures for SAFE-NET are provided in Appendix A.1. To ensure robust experimental validation, we conducted comprehensive ablation studies examining key design choices including sensitivity to the number of Fourier features (Figure 9 and 18), activation function selection for the hidden layer (Figure 17 and Table 15), sensitivity to our initialization strategy for Fourier coefficients and frequencies (Tables 17 and 18), number of layers (Figure 19), and normalization (Figure 20). These ablation results, presented Appendix C.1, explaining the choice of each component.
By combining the generality of Fourier features with known solution characteristics, SAFE-NET accelerates convergence. Figure˜3 compares SAFE-NET with the best competing PINN architecture for each PDE in Table˜2.